# A time-stepping deep gradient flow method for option pricing∑in (rough) diffusion models

### v1

@@ 6.2. Training and computational times
Table 1 summarizes the training times for the TDGF and the DGM methods in the different models. As we have expected, due to the time stepping and the absence of a second derivative in the cost function, the training of the TDGF method is faster than for the DGM method. Moreover, the training time for the TDGF method does not increase as the dimension of the problem (i.e. the number of variance processes) increases. On the contrary, the training time for the DGM method increases significantly with the number of dimensions, and becomes more than 4 times slower in the lifted Heston model when the dimension of the variance processes increases from 1 to 20.
The computational times for the TDGF and the DGM methods in all models are presented in Table 2. The computational times are taken as the average over 34 time points of computing the option prices on a grid of 41 moneyness points. Even though we take an average, there is still a high variance in the computational times. Both methods are significantly faster than the COS method in the lifted Heston model, in which we do not know the characteristic function explicitly and a system of Riccati equations needs to be solved for every evaluation of the characteristic function. The DGM method tends to be slightly faster than the TDGF method, since the latter needs to load the model corresponding to the time at which we want to compute the option price, whereas the DGM method only needs to load the weights once for all times.
Model Black–Scholes Heston lifted Heston, $n=1$ lifted Heston, $n=5$ lifted Heston, $n=20$
DGM $7.5\times 10^{3}$ $12.5\times 10^{3}$ $13.3\times 10^{3}$ $22.3\times 10^{3}$ $56.1\times 10^{3}$
TDGF $4.1\times 10^{3}$ $6.0\times 10^{3}$ $6.4\times 10^{3}$ $6.7\times 10^{3}$ $7.6\times 10^{3}$
Table 1. Training time in seconds of the different methods for a European call option in the different models.
Model Black–Scholes Heston lifted Heston, $n=1$ lifted Heston, $n=5$ lifted Heston, $n=20$
Exact/COS 0.00025 0.018 8.9 9.2 10.4
DGM 0.0043 0.0016 0.0034 0.027 10.0053
TDGF 0.024 0.0060 0.020 0.010 10.025
Table 2. Computational time in seconds of the different methods for a European call option in the different models.
@@ 7. Conclusion
We have developed in this article a novel deep learning method for option pricing in diffusion models. Starting from the option pricing PDE, using a backward differentiation time-stepping scheme and results from the calculus of variations, we can rewrite the PDE as an energy minimization problem for a suitable energy functional. The time-stepping provides a good initial guess for the next step in the optimization procedure, while the energy formulation yields an equation that only has first derivatives, instead of second ones. The energy minimization problem offers a natural candidate for a cost function, and the PDE is solved by training a neural network using stochastic gradient descent. We have considered as examples for our method the Black–Scholes model, the Heston model and the lifted Heston model, and derived the appropriate representations for the PDE in all these cases.
Overall, the TDGF method developed here performs well in the numerical experiments, with an error in the order of $10^{-3}$. Moreover, the training times for the TDGF method are faster compared to the popular DGM method, while the computational times for the lifted Heston model are much faster than the standard COS method. However, when the dimension of the variance processes in the lifted Heston model increases and the correlation increases simultaneously, then the error also increases. This increase is probably due to the hypoellipticity of the equations, and the development of suitable numerical methods will be the subject of future research.

### v2

@@ 6.2. Training and computational times
Table 6.2 summarizes the training times for the TDGF and the DGM methods in the different models. As we have expected, due to the time stepping and the absence of a second derivative in the cost function, the training of the TDGF method is faster than for the DGM method. Moreover, the training time for the TDGF method does not increase as the dimension of the problem (i.e. the number of variance processes) increases. On the contrary, the training time for the DGM method increases significantly with the number of dimensions, and becomes more than 4 times slower in the lifted Heston model when the dimension of the variance processes increases from 1 to 20.
The computational times for the TDGF and the DGM methods in all models are presented in Table 6.2. The computational times are taken as the average over 34 time points of computing the option prices on a grid of 47 moneyness points. Even though we take an average, there is still a high variance in the computational times. Both methods are significantly faster than the COS method in the lifted Heston model, in which we do not know the characteristic function explicitly and a system of Riccati equations needs to be solved for every evaluation of the characteristic function.
The possibility of splitting the computation of an option price in a slow, offline phase (training) and a very fast, online one (computation), is one of the advantages of deep learning methods such as the TDGF proposed here and the DGM. However, we should point out that when changing the model parameters, then the TDGF and the DGM need to be retrained, thus the overall time (training and computation) is slower than the evaluation of the COS method. The development of “deep parametric PDE” methods as in Glau and Wunderlich [30] would be an interesting avenue to explore in that respect. Finally, let us also mention that the TDGF requires more memory than the DGM, around 10-100MB, as it has to store a separate neural network for each time step, whereas the DGM needs only one neural network for all times; the increase in speed though justifies this cost.
{tabu}
[…]
Table 6.1. Training time in seconds of the different methods for a European call option in the different models.
{tabu}
[…]
Table 6.2. Computational time in seconds of the different methods for a European call option in the different models.
@@ 7. Conclusion
We have developed in this article a novel deep learning method for option pricing in diffusion models. Starting from the option pricing PDE, using a backward differentiation time-stepping scheme and results from the calculus of variations, we can rewrite the PDE as an energy minimization problem for a suitable energy functional. The time-stepping provides a good initial guess for the next step in the optimization procedure, while the energy formulation yields an equation that only has first derivatives, instead of second ones. The energy minimization problem offers a suitable candidate for a cost function, and the PDE is solved by training a neural network using stochastic gradient descent-type optimizers (e.g. Adam). We have considered as examples for our method the Black–Scholes model, the Heston model and the lifted Heston model, and derived the appropriate representations for the PDE in all these cases.
Overall, the TDGF method developed here performs well in the numerical experiments, with an error in the order of $10^{-3}$. Moreover, the training and computation (evaluation) times for the TDGF method are faster and more consistent compared to the popular DGM method. However, when the dimension of the variance processes in the lifted Heston model increases and the correlation increases simultaneously, then the error also increases. This increase in the error is due to the diffusion term becoming smaller, which means that the explicit term in the decomposition (3.1) becomes the dominant term. The development of suitable numerical methods for this regime will be the subject of future research.